OpenStudy (anonymous):

What percent of 24 is 18? ? %

OpenStudy (triciaal):

percent means per 100 set up a ratio 18/24 = x/100 where x is the percentage (the equivalent out of 100) solve for x to get 75%

OpenStudy (anonymous):

18/24 x 100

OpenStudy (anonymous):

so its jst 75%

OpenStudy (anonymous):

thanks everyone :)

OpenStudy (anonymous):

am not sure who i should give a medal to so confused lol

sammixboo (sammixboo):

\(\large \rm \frac{part}{whole}~=~\frac{percent}{100}\) The part is 18, since you are trying to find out what percent of 24 is 18. The whole is 24, since you are finding out the problem from this number. The percent is unknown, because that is what we are trying to find. \(\large \rm \frac{18}{24}~=~\frac{percent}{100}\) Next you can multiply the part by 100, and take the product of part x 100, and divide is by the whole. 18 x 100 = 1,800 1,800/24 = percent

sammixboo (sammixboo):

And you are correct! It's 75%

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

and love your explaination

sammixboo (sammixboo):

No problem!